Simple interest problem: How much money is in an account after 9 years and 3 months if you initially deposit $275, with a simple interest of 4.2%. Interest is added four times a year.
Answer #1
Initial deposit 275
Annual rate of interest 4.20%
Quarterly rate of interest 1.05%
Periods in quarter (9*4+1) 37
Total Simple interest = 275*1.05% *37 = 106.84
Total Amount in account after 9yrs 3months = 275 + 106.84 = 381.84